OpenStudy (anonymous):

Don't know how to solve this one.... What value of k solves the equation? k -3 = 1/27

OpenStudy (anonymous):

add 3 to both sides

OpenStudy (anonymous):

-9 is what i think it is

OpenStudy (anonymous):

or

OpenStudy (anonymous):

3

OpenStudy (anonymous):

is it \(k^{-3}=\frac{1}{27}\)

OpenStudy (anonymous):



OpenStudy (anonymous):

\[k^{-3}=\frac{1}{k^3}=\frac{1}{27}=\frac{1}{3^3}\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

So,the answer that i chose was right.

OpenStudy (anonymous):

k = 3

OpenStudy (anonymous):

Right. Thank you!

OpenStudy (anonymous):

you're welcome!